OpenStudy (mackenzie2013):

*MEDAL* Mr. Jones took a survey of college students and found that 80 out of 89 students are liberal arts majors. If a college has 10,407 students, what is the expected number of students who are liberal arts majors?

OpenStudy (anonymous):

Find the percent of the surveyed group that are liberal arts students. Then multiply that by the actual number of college students. Then subtract the number you get from the total number of college students.

OpenStudy (anonymous):

Sorry don't subtract at the end my bad.

OpenStudy (mackenzie2013):

Is there anyway you can like show the steps, Im totally confused haha

OpenStudy (texaschic101):

proportion time 80/89 = x/10,407 cross multiply 89x = (80)(10407) 89x = 832,560 x = 832,560/89 x = you can do the math

OpenStudy (anonymous):

Yea or do that. lol

OpenStudy (anonymous):

Either way you get the same answer.

OpenStudy (mackenzie2013):

so what i need yo do is take 832,560 and divide it by 89? the question also says "Round to the nearest sudent" Would this mean round?

OpenStudy (anonymous):

Round down.

OpenStudy (anonymous):

Even if it's 0.5 or over for this just round down.

OpenStudy (texaschic101):

for example : 63.2....when rounding, this = 63 or 63.5 when rounded = 64.... however, you need to round down because you can't count a number less then 1 as a whole person

OpenStudy (mackenzie2013):

I dont believe im dividing right, I keep getting 9354.60674157 as the answer to X?

OpenStudy (anonymous):

That is correct.

OpenStudy (texaschic101):

you got it right....but you need to round down...so 9354.60....rounded down = 9354

OpenStudy (anonymous):

Just round down to 9354

OpenStudy (mackenzie2013):

Agh okay, thats what confused me! Thank You!

OpenStudy (texaschic101):

YW :)

OpenStudy (anonymous):

NP